# Automated Defects Detection and Fix in Logging Statement

### v1

@@ 5.4. RQ1: Effectiveness of LogFixer
Answer to RQ1: LogFixer demonstrates strong capabilities in defect detection and updating. It significantly outperforms CodeT5+ in detection with an F1 score of 0.625. Regarding its update ability, it improves static text and dynamic variables by 48.12% and 24.90%, respectively.
5.5. RQ2: Effects of Different Components
@@
Approach. We consider the contribution of different components to LogFixer ’s performance from two perspectives.
@@
For updating capability, we compare the results with the scenario where the detection module is entirely removed and the LLM directly updates the logging statements that need to improve. This will illustrate the necessity of the detection module.
@@
Answer to RQ2: Each component of LogFixer contributes variously to its overall performance. Adjustments to hyperparameters or removal of the checker or detection components result in performance decline.
5.6. RQ3: Generalizability of LogFixer
Approach. We evaluate the capability of LogFixer using code-specific (Phind-CodeLlama (Hugging Face, 2024), Deepseek-coder (Zhu et al., 2024)) and general (Claude3.5-sonnet (Anthropic, 2024), Llama3 (AI@Meta, 2024), GPT3.5 (OpenAI., 2022), GPT4o (OpenAI, 2024)) LLMs on both synthetic data and existing logpatches. This can illustrate the generalizability of our framework, which means that we don’t rely on the ability of specific LLM backbones.
Result. Table 5 shows the performance of LogFixer with six different backbone models. First, we observe that our framework can consistently perform well on most backbones in terms of all metrics by a large margin. On average, all models get 0.614 F1-macro for detection ability. From the perspective of update ability, models have an average of IC 13.19$\uparrow$ (reflected by BLEU-4), 38.68$\uparrow$ (reflected by ROUGE-L), and 12.52$\uparrow$ (reflected by F1) respectively.
The results demonstrate the generalizability of LogFixer across various backbone models. We anticipate that the performance of LogFixer can be further enhanced with advancements in LLMs.
Table 5. The performance of LogFixer with different backbone models.
@@ 5.6. RQ3: Generalizability of LogFixer
Answer to RQ3: LogFixer showcases great generalization ability by consistently enhancing the quality of logging statements across various LLMs, including code-specific and general ones.
5.7. RQ4: Practioners’ feedback on LogFixer
In this RQ, we evaluate LogFixer based on the practitioners’ feedback from two prospective.
@@ 5.7. RQ4: Practioners’ feedback on LogFixer
Approach We aim to assess if LogFixer can accurately identify and update unseen defective logging statements while achieving a reasonable successful rate (SR). For this, we employ the new repositories dataset (Sec.  5.1.3), which has no prior identified logging statement defects and modifications. Two authors manually verified each case to confirm whether the detection is accurate and if the update is correct. If both annotators unanimously agree, we consider it a success case. Finally, we compute the success rate as the ratio of the number of successful cases to the total number of cases.
Result LogFixer reports 149 defective logging statements out of 100K data. The two authors achieve a Cohan’s Kappa of 0.84 and finally identify 91 successful detections and updates after reaching a consensus, resulting in a 61.49% SR. This outcome demonstrates the low false positive rate of LogFixer, enabling developers to conserve time in scrutinizing meaningless updates. Table 6 demonstrates SR for LogFixer across different types of defects, with the highest SR of 88.57% observed for readability issues, and the lowest at 46.15% for static-dynamic inconsistency.
@@
Approach. To understand the importance of fixing identified defects from the developers’ perspective, we randomly sample 40 updated logging cases and submit pull requests (PRs) to the repository maintainers. For each pull request, we include the defects type and the improved logging statements suggested by LogFixer. Developers’ willingness to invest effort in testing, validating, and accepting our PRs serves as a testament to the effectiveness of our approach and the perceived value of the proposed logging quality improvement.
Result. To date, 25 out of 40 proposed changes have been accepted and merged into the codebases, and an additional 3 cases have been recognized as bugs and are incorporated in future version. Table 7 shows five representative real-world defective logging statements updated by LogFixer (four successful cases and one failure case), with an in-depth discussion for each case in the “explanation” column.
Table 6. The SR for LogFixer in new data, ”SCI”, ”SDI”, ”TI”, and ”RI” indicate the statement-code, static-dynamic, temporal inconsistency, and readability issues, respectively.
Defect type SCI SDI TI RI
@@
Table 7. The real-world defective logging statement update cases suggested by LogFixer.
Information Update code Explanation
Statement-code inconsistency merged by infinispan (Anonymous, 2024b) public CompletionStage<Void> rollbackAsync (){ - log.tracef("Transaction.commit() invoked in transaction with Xid=%s", xid); + log.tracef("Transaction.rollback() invoked in transaction with Xid=%s", xid); The older statement was misleading as it implies that a commit operation was undertaken, whereas the transaction actually involves rolling back.
Static-dynamic inconsistency merged by trino (Anonymous, 2024c) private internalCompilePageWithPositionComparator (...) { - log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortChannels); + log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortOrders); The original statement logged variable sortChannels twice, mistakenly applying it to record both the channels and orders in the message.
Temporal inconsistency merged by openhab-addons (Anonymous, 2024d) private class DigiplexReceiverThread extends Thread { - logger.debug("Receiver thread started"); + logger.debug("Starting receiver thread"); Optional<String> message = readLineBlocking(); The old statement used the past tense (“started”) suggesting that the thread had already started. However, in reality, the thread initiation occurs later in the code execution.
Readability issue merged by openhab-addons (Anonymous, 2024a) public PentairIntelliFloHandler(Thing thing) { ... - logger.debug("Intellflo received refresh command"); + logger.debug("IntelliFlo received refresh command"); } The original one contains typo Intellflo. LogFixer successfully update it to IntelliFlo (a pool pump) rather than Intelliflo (a software platform).
A failure case case offered from LogFixer (false positive). public Future<Integer> scaleDown(...) { while (next > scaleTo) { next--; - log.info("Scaling down from {} to {}", next + 1, next); + log.info("Scaling down from {} to {}", next, next - 1); LogFixer erroneously reports the logging statement as defective due to a misinterpretation of next+1 as representing the current number. This error suggests that LogFixer may struggle to calculate changes in numerical variables within programs, attributing to the limited mathematical ability of LLMs.
@@
Answer to RQ4: In repositories containing previously unidentified defects, LogFixer suggests 61.49% precise logging fixes according to human evaluation, illustrating its real-world practicality. Out of 40 changes reported, developers have merged 25 cases, further validating its usefulness.
6. Threats to Validity
This section discuss the threats to validity of LogFixer.
Potential data leakage. We conduct experiments using various LLM backbones. Because the training data for these models are undisclosed, there is a risk of potential data leakage. To mitigate this, we initiate our data collection in April 2024, postdating the training of most LLMs. Furthermore, our methods also demonstrate strong performance on synthetic datasets, suggesting that the improvements in logging statements are from inference instead of memorization of training data.
Fine-tuning dataset construction. The fine-tuning dataset construction process may raise concerns. Firstly, similar to logging-related works (Mastropaolo et al., 2022; Li et al., 2020; Liu et al., 2022), the basic assumption is that the logging statements in the latest programs are defect-free. Secondly, due to the required fine-tuning set scale, manually labeling a large number of negative examples would be impractical. Consequently, we employed mutation techniques to automatically synthesize defective examples, potentially leading to the inclusion of suboptimal negative examples. To mitigate this, the selection of data source is limited to well-maintained codebases (details in 4.1), which are considered as positive data. Additionally, we devised various mutation methods to ensure the diversity, realism, and scalability of synthetic data.
The selection of programming languages. We source data from 641 Java projects, which may impact the generalizability of our method for other languages. Following previous study (Ding et al., 2023), we choose Java because it is a mature language with multiple built-in logging and third-party logging frameworks. Nevertheless, applying LogFixer to other programming languages may affect its performance because of distinct language features. While fine-tuning detectors on other language-specific datasets is a promising solution, we aim to incorporate more languages in future work.

### v2

@@ 5.3. RQ1: Effectiveness of LogUpdater
Answer to RQ1: LogUpdater demonstrates strong capabilities in defect detection and updating. It outperforms all baselines in detection with an F1 score of 0.625. Regarding its update ability, it improves static text and dynamic variables by 48.12% and 24.90%, respectively.
5.4. RQ2: The end-to-end effectiveness of LogUpdater
Approach. In this RQ, our goal is to explore the efficacy of LogUpdater once it integrates the tasks of detecting and updating defective logging statement. We conduct a thorough evaluation and comparison between LogUpdater and several leading LLMs: a general-purpose model (GPT4o), a code-based model (Deepseek-coder-v2), and a reasoning model (Deepseek-R1). For each of these LLM-based benchmark models, we supply five examples for ICL to facilitate task demonstration. Details of the prompt used for the end-to-end task are available in  (Logupdater, 2025a). We assess these models using the existing logpatches dataset, as elaborated in Section 5.1.1, focusing on metrics such as Detection Precision, Recall, F1-Score, ROUGE-L, and Variable-F1. It should be noted that for each model, the computation of updating ability metrics is limited to cases in which a defective detection is successfully achieved. Additionally, to appraise the financial cost associated with each method, we introduce the concept of cost per input (CPI), representing the average cost of using the official API of LLM for each input that aims at the detection and revision of defective logging statements. Note that the base LLM model for LogUpdater is GPT4o.
Result. Table 6 presents the evaluation results for end-to-end comparisons of LogUpdater with the baselines. We can see the LogUpdater outperforms all baselines in terms of all metrics. LogUpdater demonstrates significant performance enhancements compared to all baseline models when evaluated using the metrics of Precision, Recall, and F1-score. Specifically, LogUpdater achieves a precision of 0.725 and recall of 0.552, resulting in an F1-score of 0.625, which significantly surpasses the baseline models (0.373, 0.372, and 0.337 for GPT4o, Deepseek-coder-v2, and Deepseek-R1, respectively). That means, for the detection performance, LogUpdater detects more defectifve logging statements with higher precision.
In terms of the updating metrics, we observe that LogUpdater demonstrates competitive performance in ROUGE-L (0.928) and Variable-F1 (0.977) metrics. Notably, LogUpdater shows substantial improvements over original logging statements as indicated by the percentage increases: 81.29% in ROUGE-L and 55.77% in Variable-F1. This suggests that LogUpdater is particularly effective at fixing the defective logging statements at both the static content aspect and dynamic variable aspect.
From an efficiency perspective, LogUpdater demonstrates remarkable cost-effectiveness compared to its base model GPT4o. With a CPI of 0.0011, LogUpdater is approximately 16.2 times more cost-efficient than the direct use of GPT4o (0.0194). This substantial cost reduction while maintaining superior performance is particularly noteworthy since both approaches utilize GPT4o as their foundation. The significant cost advantage of LogUpdater can be attributed to its specialized architecture (i.e., use small model to develop a detector), which minimizes the number of API calls required for each input cases analysis. This cost-effectiveness, combined with LogUpdater’s superior detection and generation performance, makes it a more viable solution for large-scale logging statement maintenance in real-world applications.
Table 6. The end-to-end comparisons of LogUpdater with the baselines.
Precision Recall F1-score ROUGE-L Variable-F1 CPI
GPT4o 0.401 0.403 0.373 0.905 (68.33$\uparrow$) 0.967 (29.79$\uparrow$) 0.0194
Deepseek-coder-v2 0.383 0.365 0.372 0.861 (51.90$\uparrow$) 0.979 (0.00%) 0.0022
Deepseek-R1 0.358 0.414 0.337 0.897 (62.68$\uparrow$) 0.905 (12.03$\uparrow$) 0.0044
LogUpdater 0.725 0.552 0.625 0.928 (81.29$\uparrow$) 0.977 (55.77$\uparrow$) 0.0011
Answer to RQ2: In summary, LogUpdater significantly outperforms baseline models in terms of all metrics by significant margins. Moreover, LogUpdater is 16.2 times more cost-efficient than GPT4o, enhancing its viability for practical applications.
5.5. RQ3: Effects of Different Components
@@
Approach. We consider the contribution of different components to LogUpdater ’s performance from two perspectives.
@@
To assess the updating capability, we begin by comparing the outcomes with a scenario in which the detection module is completely omitted, allowing the LLM to directly update the logging statements requiring enhancement. This approach will highlight the importance of including the detection module. Subsequently, we conduct an analysis to evaluate how varying the number of similar LCCs presented in the prompt influences the effectiveness of the Updater.
@@
Table 7 demonstrates the performance during updating across varying numbers of similar LCCs as outlined in the prompt. It is evident that all performance metrics show a gradual increase when the number of shots is incremented from 1 to 4, for the case of existing log patches, and up to 3 for synthetic data. This trend reaches its highest point at 4 shots, recording a peak value of 0.395 for BLEU-4, which signifies a 2.8% increase compared to the 1-shot scenario. Such findings indicate that the updater gains from a more substantial provision of examples in the prompt, thereby facilitating an improved contextual comprehension and consequently enhancing the preciseness of the logging statements. Nonetheless, beyond this optimal range, when the number of shots is further escalated to 5 and 10, a slight decrease is observed across all metrics.
@@
Table 7. Shot Number Analysis
Shot Number 1 2 3 4 5 10
Existing logpatches
BLEU-4 0.384 0.387 0.388 0.395 0.391 0.385
ROUGE-L 0.848 0.850 0.855 0.864 0.858 0.844
Variable-F1 0.445 0.438 0.447 0.462 0.449 0.434
Synthetic data
BLEU-4 0.436 0.440 0.446 0.443 0.442 0.438
ROUGE-L 0.751 0.757 0.762 0.759 0.760 0.754
Variable-F1 0.787 0.792 0.792 0.796 0.789 0.784
Answer to RQ3: Each component of LogUpdater contributes variously to its overall performance. Removing the detection module notably degrades performance, underscoring its critical role in enhancing logging statement accuracy. 4-shot LCCs in prompt maximizes the update performance of LogUpdater.
5.6. RQ4: Generalizability of LogUpdater
Approach. We evaluate the capability of LogUpdater using code-specific (Phind-CodeLlama (Hugging Face, 2024), Deepseek-coder (Zhu et al., 2024)) and general (Claude3.5-sonnet (Anthropic, 2024), Llama3 (AI@Meta, 2024), GPT3.5 (OpenAI., 2022), GPT4o (OpenAI, 2024)) LLMs on both synthetic data and existing logpatches. This can illustrate the generalizability of our framework, which means that we don’t rely on the ability of specific LLM backbones.
Result. Table 8 shows the performance of LogUpdater with six different backbone models. First, we observe that our framework can consistently perform well on most backbones in terms of all metrics by a large margin. On average, all models get 0.614 F1-macro for detection ability. From the perspective of update ability, models have an average of IC 13.19$\uparrow$ (reflected by BLEU-4), 38.68$\uparrow$ (reflected by ROUGE-L), and 12.52$\uparrow$ (reflected by F1) respectively.
The results demonstrate the generalizability of LogUpdater across various backbone models. We anticipate that the performance of LogUpdater can be further enhanced with advancements in LLMs.
Table 8. The performance of LogUpdater with different backbone models.
@@ 5.6. RQ4: Generalizability of LogUpdater
Answer to RQ4: LogUpdater showcases great generalization ability by consistently enhancing the logging statements across various LLMs, including code-specific and general ones.
5.7. RQ5: Practitioners’ feedback on LogUpdater
In this RQ, we evaluate LogUpdater based on the practitioners’ feedback from two perspectives.
@@ 5.7. RQ5: Practitioners’ feedback on LogUpdater
Approach. We aim to assess if LogUpdater can accurately identify and update unseen defective logging statements while achieving a reasonable successful rate (SR). For this, we employ the new repositories dataset (Sec.  5.1.3), which has no prior identified logging statement defects and modifications. Two authors manually verified each case to confirm whether the detection is accurate and if the update is correct. If both annotators unanimously agree, we consider it a success case. Finally, we compute the success rate as the ratio of the number of successful cases to the total number of cases.
Result. LogUpdater reports 149 defective logging statements out of 100K data. The two authors achieve a Cohen’s Kappa of 0.84 and finally identify 91 successful detections and updates after reaching a consensus, resulting in a 61.49% SR. This outcome demonstrates the low false positive rate of LogUpdater, enabling developers to conserve time in scrutinizing meaningless updates. Table 9 demonstrates SR for LogUpdater across different types of defects, with the highest SR of 88.57% observed for readability issues, and the lowest at 46.15% for static-dynamic inconsistency.
@@
Approach. In order to comprehensively grasp the significance of rectifying identified defects from the viewpoint of developers, we conduct a randomized sampling of 40 cases from the previously 91 instances of successfully updated logging cases. We then submit pull requests (PRs) to the repository maintainers for each sampled instance. Each pull request was accompanied by detailed information regarding the types of defects encountered, in addition to the enhanced logging statements suggested by LogUpdater. The developers’ readiness and commitment to dedicate their efforts towards testing, validating, and ultimately accepting our submitted PRs underscores the effectiveness and practical value of our proposed strategy. The full merged cases are shown in our replication package (Logupdater, 2024).
Result. To date, 25 out of 40 proposed PRs have been accepted and merged into the codebases, and an additional 3 cases have been recognized as bugs and are incorporated in future versions.
Table 9. The successful rate for LogUpdater in new data.
Defect type statement-code incon. static-dynamic incon. temporal incon. readability issues
@@
Table 10. The real-world defective logging statement update cases suggested by LogUpdater.
Information Update code
Statement-code inconsistency merged by infinispan (LogUpdater, 2024a) public CompletionStage<Void> rollbackAsync (){ - log.tracef("Transaction.commit() invoked with Xid=%s", xid); + log.tracef("Transaction.rollback() invoked with Xid=%s", xid); status = Status.STATUS_MARKED_ROLLBACK;
Static-dynamic inconsistency merged by trino (LogUpdater, 2024b) private Comparator internalCompilePageWithPositionComparator(List<Type> types, List<Integer> sortChannels, List<SortOrder> sortOrders){ - log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortChannels); + log.error(t, "Error compiling comparator for channels %s with order %s", sortChannels, sortOrders); comparator = new SimplePageWithPositionComparator(types, sortChannels, sortOrders, typeOperators);
Temporal inconsistency merged by openhab-addons (LogUpdater, 2024c) public void run() { - logger.debug("Receiver thread started"); + logger.debug("Starting receiver thread"); while (!interrupted()) { Optional<String> message = readLineBlocking();} logger.debug("Receiver thread finished");
Readability issue merged by openhab-addons (LogUpdater, 2024) public void handleCommand(ChannelUID cUID, Command cmd) { if (command instanceof RefreshType) { logger.debug("Intellflo received refresh command"); logger.debug("IntelliFlo received refresh command"); updateChannel(channelUID.getId(), null);}}
A false positive case offered from LogUpdater. public Future scaleDown(Reconciliation reconciliation, String namespace, String name, int scaleTo, long timeoutMs) { Integer nextReplicas = currentScale(namespace, name); while (nextReplicas > scaleTo) { nextReplicas--; LOGGER.infoCr(reconciliation, "Scaling down from {} to {}", nextReplicas + 1, nextReplicas); resource(namespace, name).withTimeoutInMillis(timeoutMs).scale(nextReplicas);} return nextReplicas;}
@@
Answer to RQ5: In repositories containing previously unidentified defects, LogUpdater suggests 61.49% precise logging fixes according to human evaluation, illustrating its real-world practicality. While this demonstrates considerable effectiveness on unseen data, it also highlights that expert validation remains an essential part of the workflow, although for these successful cases, the effort is largely reduced to reviewing the proposed fix. Furthermore, out of 40 changes reported, developers have merged 25 cases, further validating its usefulness.
5.8. Case Study
Table 10 shows five representative real-world defective logging statements updated by LogUpdater (four successful cases and one failure case). In the first four success cases: (1) The statement-code inconsistency merged by infinispan (LogUpdater, 2024a) fixes the misleading as the older logging statement implies that a commit operation was undertaken, whereas the transaction actually involves rolling back. (2) In the static-dynamic inconsistency case merged by trino (LogUpdater, 2024b), the original statement logged variable sortChannels twice, mistakenly applying it to record both the channels and orders in the message. (3) For the temporal inconsistency case merged by openhab-addons (LogUpdater, 2024c), The old statement used the past tense (“started”), suggesting that the thread had already started. However, in reality, the thread initiation occurs later in the code execution. (4) In the readability issue merged by openhab-addons (LogUpdater, 2024), the original one contains a typo Intellflo. LogUpdater successfully update it to IntelliFlo (a pool pump) rather than Intelliflo (a software platform). These real-world updates, which were confirmed by developers, indicate the practicality of LogUpdater in detecting and fixing various defective logging statements.
The last row shows a failure (false positive) case offered by LogUpdater, where it erroneously reports the logging statement as defective due to a misinterpretation of next+1 as the current number. This error highlights LogUpdater’s difficulty in handling numerical variable changes, likely due to the limited mathematical capabilities of LLMs. Despite several false positive cases, LogUpdater has significantly reduced the manual effort required to inspect all logging statements in a project, enhancing development efficiency and reliability. To eliminate these issues in the future direction, an advanced program logical comprehension module could be introduced.
6. Discussion
6.1. Advantages of LogUpdater
The effectiveness of LogUpdater has been evidenced by the new project dataset, with 25 out of 40 suggested modifications being accepted and integrated into widely-used, real-world repositories. We expand upon LogUpdater in the context of actual logging practices.
6.1.1. Benefits of utilizing Checker component
The Checker is designed to mitigate false positives from the similarity-based detector, which are prevalent due to the highly imbalanced nature of real-world data (i.e., most logging statements are not defective). By comparing the effectiveness of Checker in both balanced data and unbalanced data, we argue that Checker is quite beneficial for the actual development scenario. For the balanced dataset, Figure 5 shows a noticeable improvement in precision (increases of 0.15 and 0.11) with a modest reduction in recall, confirming its role in reducing false positives as intended. The value of the Checker emerges in unbalanced datasets. For example, in our new project dataset with 100K entries, the similarity-based classifier initially identified 1,771 potential defective logging statements, and only 149 of them remained after passing Checker. We further manually checked 100 randomly selected cases considered false positives by the Checker and revealed only 4 false negatives, demonstrating a 96% filtering accuracy while saving the cost of LLM invocation in the subsequent step.
6.1.2. Applicability of automatically generated logging statements
Although recent works have investigated the capabilities of LLMs in generating logging statements automatically (Li et al., 2023b, 2024a), LogUpdater continues to offer unique benefits from the following two aspects. First, the generated logging statements is far from mature. LLMs exhibit potential limitations, particularly in accurately generating static text intertwined with dynamic variables — a key area prone to defects (Li et al., 2023b). For example, the logging statement generated by the SOTA solution (Li et al., 2024a) achieved a ROUGE-L score of 0.509. However, the ROUGE-L score comparing the logging statements before and after commit changes was 0.707. The performance of LLMs in automatically writing logging statements is still far from being applicable, let alone enabling them to avoid various complex defects. Our LogUpdater is designed to address these intricacies, identifying and rectifying such defects. Second, while LLM-based generation approaches are effective in generating new logging statements to some extent, they are not typically utilized to retrospectively analyze and update large amounts of existing logging statements within codebases due to computational and practical constraints. LogUpdater fills this gap by efficiently processing vast quantities of existing logging statements, ensuring they meet current standards without significant overhead.
6.1.3. Capability to solve multiple defects in one logging statement situation
During our manual analysis, we find one defective logging statement case (1/258, 0.3%) containing multiple defects—specifically, a combination of statement-code inconsistency and readability issues. This finding indicates that it is indeed quite rare for a logging statement to simultaneously exhibit multiple types of defects. Our approach is capable of addressing related issues. The detection model still provides classification results for these special cases, following which Updater performs updates based on the corresponding prompt.
6.1.4. Capability for extending new defect categories
In this study, we conducted a manual analysis of log-centric commits sampled from 641 repositories, identifying four distinct defect types, which collectively account for 51.6%—over half of the cases. It is possible that specific repositories may exhibit unique new defect types. Our proposed method is well-suited to adapt to such scenarios. By retraining the similarity-based classifier and making minor adjustments to the prompt, our approach can seamlessly extend to incorporate one or more additional defect categories.
6.1.5. Cost-management for LogUpdater’s pipeline design
LogUpdater’s architecture prioritizes cost-effectiveness, particularly in environments with numerous logging statements. By separating the detection and updating phases, LogUpdater efficiently filters out non-defective logging statements, minimizing the need for expensive LLM invocations. With a CPI of 0.0011, LogUpdater is approximately 16.2 times more cost-efficient than the baseline model GPT4o (CPI of 0.0194). This significant cost reduction, achieved while maintaining superior performance metrics, is primarily due to fewer API calls required for processing. In practical scenarios, the potential savings could exceed those measured in RQ2, as most logging statements do not need LLM processing. By focusing only on relevant context, LogUpdater reduces the information sent to the LLM, enhancing cost-efficiency. This makes LogUpdater a compelling solution for practitioners seeking to improve logging practices without incurring high costs.
6.2. Threats to Validity
6.2.1. Construct validity
Fine-tuning dataset construction. The fine-tuning dataset construction process may raise concerns. Firstly, similar to logging-related works (Mastropaolo et al., 2022; Li et al., 2020; Liu et al., 2022), the basic assumption is that the logging statements in the latest programs are defect-free. Secondly, due to the required fine-tuning set scale, manually labeling a large number of negative examples would be impractical. Consequently, we employed mutation techniques to automatically synthesize defective examples, potentially leading to the inclusion of suboptimal negative examples. To mitigate this, the selection of data source is limited to well-maintained codebases (details in 4.1), which are considered as positive data. Additionally, we devised various mutation methods to ensure the diversity, realism, and scalability of synthetic data. Limitation in scope of defect types. While our work focuses on four factually verifiable defect types (e.g., semantic inconsistencies), we intentionally excluded logging-level changes (24.8%) and other subjective practices. This exclusion introduces a threat to the method’s comprehensiveness, as misuse logging-level issues are critical to system observability. However, our decision aligns with the prioritization of objective, context-agnostic defects to build a robust automated framework. To mitigate this threat, we plan to enhance our framework by supporting context-aware misuse log-level detection in the future work. Repair Effectiveness Measurement. A potential concern pertains to the practical implication of the 61.49% success rate reported in RQ5. We acknowledge that this rate, while demonstrating the tool’s capability, signifies that manual intervention by domain experts is still necessary to validate the suggestions and correct the remaining 38.51% of cases where the detection or the proposed fix might be inaccurate. It is important to clarify that LogUpdater is designed to significantly reduce and shift the manual workload associated with logging statement maintenance, rather than eliminating it entirely. Compared to fully manual inspection or approaches relying solely on defect detection tools (which necessitate manual diagnosis and fixing for 100% of identified potential issues), LogUpdater offers substantial effort savings. Firstly, for the 61.49% successful cases, the developer effort is reduced from diagnosis and correction to primarily validation of the automatically generated fix. Secondly, even for the unsuccessful suggestions, the tool focuses developer attention on specific problematic statements and provides a starting point for correction. Therefore, while human oversight remains crucial, LogUpdater represents a significant step towards automating log quality improvement by handling a substantial portion of the detection and repair tasks, thereby reducing the overall manual burden.
6.2.2. Internal validity